# DiEC: Diffusion Embedded Clustering

### v1

@@ Abstract
Deep clustering hinges on learning representations that are inherently clusterable. However, using a single encoder to produce a fixed embedding ignores the representation trajectory formed by a pretrained diffusion model across network hierarchies and noise timesteps, where clusterability varies substantially. We propose DiEC (Diffusion Embedded Clustering), which performs unsupervised clustering by directly reading internal activations from a pretrained diffusion U-Net.
DiEC formulates representation selection as a two-dimensional search over layer $\times$ timestep, and exploits a weak-coupling property to decompose it into two stages. Specifically, we first fix the U-Net bottleneck layer as the Clustering Middle Layer (CML), and then use Optimal Timestep Search (OTS) to identify the clustering-optimal timestep ($t^{*}$). During training, we extract bottleneck features at the fixed $t^{*}$ and obtain clustering representations via a lightweight residual mapping. We optimize a DEC-style KL self-training objective, augmented with adaptive graph regularization and entropy regularization to strengthen cluster structures. In parallel, we introduce a denoising-consistency branch at random timesteps to stabilize the representations and preserve generative consistency. Experiments show that DiEC achieves competitive clustering performance on multiple standard benchmarks.
@@ 1 Introduction
Figure 1: Overview of DiEC (Diffusion Embedded Clustering). Given an input image $x_{0}$, we inject noise using both a random timestep $t_{\mathrm{rand}}$ and a fixed clustering-optimal timestep (COT). The random-timestep branch preserves generative consistency via standard denoising, while the fixed-timestep branch extracts the Clustering Middle Layer(CML) for clustering-aware optimization. We apply residual decoupling on CML and jointly optimize a graph regularizer and a KL-based clustering objective to obtain more stable and discriminative clustering representations.
@@
Clustering is a fundamental task in unsupervised learning, aiming to uncover latent class structure from unlabeled data, and has been widely used in applications such as image understanding vangansbeke2020scan, anomaly detection pang2021deepanomalyreview, and large-scale retrieval johnson2019billion. However, modern data are often high-dimensional, nonlinear, and corrupted by noise, making clustering increasingly dependent on high-quality and robust feature representations zhou2024csurdeepclustering. Traditional approaches typically rely on shallow features or simple distance metrics, and thus struggle to capture deep semantic structure in complex data bendavid2018clusteringwrong. Recent studies have therefore introduced deep models into clustering by jointly optimizing representation learning and clustering objectives, improving performance in challenging settings zhou2024csurdeepclustering. This line of research is commonly referred to as deep clustering and has been validated by representative methods such as Ren et al. ren2024tnnlsdeepclustering and Wei et al. wei2024overviewdeepclustering.
@@
Recent progress in deep clustering under high-dimensional, nonlinear, and noisy settings has been largely driven by learning more semantic representations with deep networks, which in turn induces clearer cluster structures in the learned space zhou2024survey. Accordingly, much of the literature centers on the question of how to obtain clustering-friendly representations zhou2024survey. Existing approaches have broadly evolved along two complementary lines: (i) generative encoders, represented by autoencoders, variational autoencoders, and GANs, which learn low-dimensional embeddings via reconstruction or generation objectives xie2016dec; jiang2017vade; mukherjee2019clustergan; and (ii) discriminative self-supervised learning, represented by contrastive learning, which enforces consistency constraints to learn discriminative representations vangansbeke2020scan; caron2020swav. Despite the different training signals, both families typically share a common modeling assumption: the feature extractor naturally yields a single, most clustering-discriminative primary representation, on top of which cluster assignment and structure refinement are performed xie2016dec; zhou2024survey.
@@
We observe that the clusterability of internal representations in diffusion denoising networks is not random; instead, it varies systematically with both the noise scale and the network hierarchy. Along the noise-scale dimension, clustering performance typically follows a unimodal trend, improving at moderate noise levels and degrading as the noise becomes too strong. Along the hierarchy dimension, representations around the bottleneck strike a more favorable balance between stability and separability. Moreover, across datasets, bottleneck activations are generally easier to access and more stable, whereas representations after the bottleneck often incur higher extraction cost with limited gains, resulting in a lower overall cost–benefit trade-off.
@@
Motivated by the above patterns, we propose Diffusion Embedded Clustering (DiEC), a framework that performs unsupervised clustering by leveraging internal embeddings from the denoising network of a pretrained diffusion model. DiEC treats internal activations generated under different noise conditions and network hierarchies as a searchable representation trajectory, and builds the clustering pipeline on top of it: we select bottleneck activations as the Clustering Middle Layer (CML) and efficiently identify the clustering-optimal timestep (COT) via a fast scan on a data subset. We then introduce clustering optimization in this representation space, sharpening assignments with a $\mathrm{KL}$-divergence-based clustering loss, preserving local structure through adaptive graph embedding, and alleviating potential conflicts between denoising and clustering objectives via residual feature decoupling, which explicitly strengthens cluster structure and improves training stability. Experiments and systematic ablations show that DiEC outperforms existing deep clustering methods on multiple benchmark datasets.Our contributions are as follows:
@@
  • We introduce a representation-trajectory perspective and show that clusterability varies systematically over the timestep $\times$ layer space; based on this insight, we identify a more clustering-friendly layer location and develop an efficient procedure to quickly localize the clustering-optimal timestep.
@@ 2 Related Work
In recent years, a growing body of work has explored using diffusion models as feature extractors to provide discriminative representations for downstream vision tasks; for a systematic overview, we refer to Fuest et al. fuest2024survey. Existing studies that leverage diffusion models for representation extraction can be broadly categorized into two groups.
@@ 3 Background
Diffusion models.Diffusion models have recently emerged as a leading paradigm for high-fidelity generative modeling. Compared with traditional generative models such as VAEs and GANs, they often offer a better trade-off between sample fidelity and training stability, and have achieved strong results in image synthesis, text-to-image generation, image editing, super-resolution, and conditional generation dhariwal2021diffusion; nichol2022glide; saharia2021sr3; saharia2022imagen.
The core idea of diffusion models is to cast generation as an iterative denoising process: they first define a forward diffusion process that gradually perturbs data into approximately isotropic Gaussian noise, and then learn a corresponding reverse denoising process that progressively transforms noise back to the data distribution.
In DDPM ho2020ddpm, given a data sample $x_{0}\sim q_{\mathrm{data}}(x)$, where $q_{\mathrm{data}}(x)$ denotes the true data distribution, the forward diffusion process is defined as a Markov chain:
$q(x_{t}\mid x_{t-1})=\mathcal{N}\!\left(x_{t};\,\sqrt{1-\beta_{t}}\,x_{t-1},\,\beta_{t}\mathbf{I}\right).$ (1)
where $\{\beta_{t}\}_{t=1}^{T}$ denotes the noise schedule and $t\in\{1,\ldots,T\}$ indexes the diffusion timestep. Let $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. Then,
@@
indicating that the noise level of a sample is governed by the timestep $t$.
The reverse process is modeled by a learnable conditional distribution:
@@
where
@@
This objective can be viewed as a simplified form of the evidence lower bound (ELBO). At sampling time, starting from $x_{T}\sim\mathcal{N}(0,\mathbf{I})$, we iteratively apply $p_{\theta}(x_{t-1}\mid x_{t})$ to progressively denoise and obtain generated samples.
@@ 4 Diffusion Embedded clustering
We denote by $x_{t,i}$ the observation of the $i$-th sample at diffusion timestep $t$, and by $x_{0,i}$ the corresponding clean input. Given an unlabeled dataset $\mathcal{D}=\{x_{0,i}\}_{i=1}^{N}$ and the number of clusters $K$, DiEC aims to learn clustering representations from intermediate activations of a pretrained diffusion model and to predict a cluster assignment $\hat{y}_{i}\in\{1,\ldots,K\}$ for each sample. DiEC first locates a clustering-friendly diffusion representation at a specific layer–timestep pair $(l^{*},t^{*})$. Conditioning on this location, it then jointly optimizes the model with a clustering objective and a structural regularizer, yielding stable and separable clustering embeddings. Figure 1 provides an overview of the framework, and additional details are presented in the following subsections.
@@ 4.2 Clustering-Aware Optimization
Clustering objective. For each sample $x_{0,i}$, we apply forward noising at the fixed optimal timestep $t^{*}$ to obtain a noisy observation
$x_{t^{*},i}=\sqrt{\bar{\alpha}_{t^{*}}}\,x_{0,i}+\sqrt{1-\bar{\alpha}_{t^{*}}}\,\epsilon,\qquad\epsilon\sim\mathcal{N}(0,\mathbf{I}).$ (19)
To smooth the stochasticity introduced by noise, we independently sample $M$ noise realizations at $t^{*}$ for the same sample, yielding $\{x_{t^{*},i}^{(m)}\}_{m=1}^{M}$, and extract the corresponding representations $\{z_{i}^{(m)}\}_{m=1}^{M}$ via the residual module. We then average them to obtain a stabilized feature
@@
We initialize cluster centroids $\{\mu_{k}\}_{k=1}^{K}$ using $k$-means on $\{z_{i}\}_{i=1}^{N}$ and then adopt a DEC-style self-training procedure: given the current centroids, we compute the soft assignments $Q=[q_{ik}]$ and construct the target distribution $P=[p_{ik}]$, and minimize
$\mathcal{L}_{\mathrm{KL}}=\mathrm{KL}(P\|Q)=\sum_{i}\sum_{k}p_{ik}\log\frac{p_{ik}}{q_{ik}},$ (21)
which progressively sharpens high-confidence assignments and strengthens cluster structure.
@@ 5 Experiment
Datasets. We evaluate DiEC on four widely used benchmarks for unsupervised clustering: MNIST, USPS, Fashion-MNIST, and CIFAR-10. MNIST and Fashion-MNIST each contain 70,000 grayscale images of resolution $28\times 28$. USPS contains 9,298 grayscale handwritten-digit images with an original resolution of $16\times 16$. CIFAR-10 consists of 60,000 color images of resolution $32\times 32$. All datasets have $10$ ground-truth classes. In our clustering setup, we treat all samples as unlabeled during training and set $K$ to the true number of classes only for evaluation when computing clustering metrics. To match the diffusion model input, all images are normalized as a preprocessing step.

### v2

@@ Abstract
Deep clustering critically depends on representations that expose clear cluster structure, yet most prior methods learn a single embedding with an autoencoder or a self-supervised encoder and treat it as the primary representation for clustering. In contrast, a pretrained diffusion model induces a rich representation trajectory over network layers and noise timesteps, along which clusterability varies substantially. We propose Diffusion Embedded Clustering (DiEC), an unsupervised clustering framework that exploits this trajectory by directly leveraging intermediate activations of a pretrained diffusion U-Net. DiEC formulates representation selection over layer $\times$ timestep and adopts a practical two-stage procedure: it uses the U-Net bottleneck as the Clustering Middle Layer (CML; $l^{*}$) and identifies the Clustering-Optimal Timestep (COT; $t^{*}$) via an efficient subset-based, noise-averaged search. Conditioning on $(l^{*},t^{*})$, DiEC learns clustering embeddings through a lightweight residual mapping, optimized with a DEC-style KL self-training objective and structural regularization, while a parallel random-timestep denoising-consistency loss stabilizes training and preserves diffusion behavior. Experiments on standard benchmarks demonstrate that DiEC achieves strong clustering performance and reveal the importance of selecting diffusion representations for clustering.
@@ 1 Introduction
Figure 1: Overview of DiEC (Diffusion Embedded Clustering). Given an input image $x_{0}$, we inject noise using both a random timestep $t_{\mathrm{rand}}$ and a fixed clustering-optimal timestep (COT). The random-timestep branch preserves generative consistency via standard denoising, while the fixed-timestep branch extracts the Clustering Middle Layer (CML) for clustering-aware optimization. We apply residual decoupling on CML and jointly optimize a graph regularizer and a KL-based clustering objective to achieve more stable and discriminative clustering representations.
@@
Clustering is a fundamental task in unsupervised learning, aiming to uncover latent class structure from unlabeled data, and has been widely used in applications such as image understanding vangansbeke2020scan, anomaly detection pang2021deepanomalyreview, and large-scale retrieval johnson2019billion. However, modern data are often high-dimensional, nonlinear, and corrupted by noise, making clustering increasingly dependent on high-quality and robust feature representations zhou2024csurdeepclustering. Traditional approaches typically rely on shallow features or simple distance metrics, and thus struggle to capture deep semantic structure in complex data bendavid2018clusteringwrong. Recent studies have therefore introduced deep models into clustering by jointly optimizing representation learning and clustering objectives, improving performance in challenging settings zhou2024csurdeepclustering. This line of research is commonly referred to as deep clustering and has been validated by representative methods such as those of ren2024tnnlsdeepclustering and wei2024overviewdeepclustering.
@@
Existing deep clustering methods have made progress in high-dimensional, nonlinear, and noisy settings largely because deep networks can learn more semantically meaningful representations, thereby improving cluster separability in the resulting embedding space zhou2024survey. Accordingly, much of the literature centers on how to obtain clustering-friendly representations zhou2024survey. Existing approaches have broadly evolved along two complementary lines: (i) generative encoders, represented by autoencoders, variational autoencoders, and GANs, which learn low-dimensional embeddings via reconstruction or generation objectives xie2016dec; jiang2017vade; mukherjee2019clustergan; and (ii) discriminative self-supervised learning, represented by contrastive learning, which enforces consistency constraints to learn discriminative representations vangansbeke2020scan; caron2020swav. Despite the different training signals, both families typically share a common modeling assumption: the feature extractor naturally yields a single, most clustering-discriminative primary representation, on top of which cluster assignment and structure refinement are performed xie2016dec; zhou2024survey.
@@
We observe that the clusterability of internal representations in diffusion denoising networks is not random; instead, it varies systematically with both the noise scale and the network hierarchy. Along the noise scale, clustering performance often follows a unimodal trend, improving at moderate noise levels and degrading at high noise levels. Along the network hierarchy, clusterability typically peaks in the early upsampling stages immediately after the bottleneck, where representations remain stable while becoming increasingly separable. Moreover, across datasets, bottleneck activations are easy to extract and provide a reliably strong clustering signal, though they are not necessarily optimal. In contrast, post-bottleneck representations often require higher extraction cost with only marginal gains, which motivates us to focus on bottleneck features in our design.
@@
Motivated by the above patterns, we propose Diffusion Embedded Clustering (DiEC), a framework that performs unsupervised clustering by leveraging internal representations from the denoising network of a pretrained diffusion model. DiEC treats internal activations generated under different noise conditions and network hierarchies as a searchable representation trajectory, and builds the clustering pipeline on top of it: we select the bottleneck layer as the Clustering Middle Layer (CML) and efficiently identify the Clustering-Optimal Timestep (COT) via a fast scan on a data subset. We then introduce clustering optimization in this representation space, sharpening assignments with a $\mathrm{KL}$-divergence-based clustering loss, preserving local structure through adaptive graph embedding, and alleviating potential conflicts between denoising and clustering objectives via residual feature decoupling, which strengthens cluster structure and improves clustering performance. Experiments and systematic ablations show that DiEC outperforms existing deep clustering methods on multiple benchmark datasets. Our contributions are as follows:
@@
  • We introduce representation-trajectory perspective and show that clusterability varies systematically over the timestep $\times$ layer space; based on this insight, we identify a more clustering-friendly layer location and develop an efficient procedure to quickly localize the clustering-optimal timestep.
@@ 2 Related Work
In recent years, a growing body of work has explored using diffusion models as feature extractors to provide discriminative representations for downstream vision tasks; for a systematic overview, we refer to fuest2024survey for a comprehensive survey. Existing studies that leverage diffusion models for representation extraction can be broadly categorized into two groups.
@@ 3 Background
Diffusion models. Compared with traditional generative models such as VAEs and GANs, diffusion models often offer a better trade-off between sample fidelity and training stability, and have demonstrated strong performance in image synthesis, text-to-image generation, image editing, super-resolution, and other conditional generation tasks dhariwal2021diffusion; nichol2022glide; saharia2021sr3; saharia2022imagen. These advances have established diffusion models as a leading paradigm for high-fidelity generative modeling.
The core idea of diffusion models is to formulate generation as iterative denoising: a forward process gradually corrupts data with Gaussian noise, and a learned reverse-time process progressively removes noise to generate samples from the data distribution.
In DDPM ho2020ddpm, the forward diffusion process is defined as a Markov chain that gradually corrupts a data sample $x_{0}\sim q_{\mathrm{data}}(x)$:
$q(x_{t}\mid x_{t-1})=\mathcal{N}\!\left(x_{t};\,\sqrt{1-\beta_{t}}\,x_{t-1},\,\beta_{t}\mathbf{I}\right),$ (1)
where $\{\beta_{t}\}_{t=1}^{T}$ is a predefined noise schedule and $t\in\{1,\ldots,T\}$ denotes the diffusion timestep. Let $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. Then, the marginal distribution of $x_{t}$ admits a closed-form expression
@@
which shows that the noise level of $x_{t}$ is fully governed by the timestep $t$.
The reverse-time process is parameterized as a Gaussian transition
@@
where the mean is expressed via a noise-prediction network $\epsilon_{\theta}$ as
@@
This objective can be interpreted as a simplified variational bound on the data likelihood. At sampling time, starting from $x_{T}\sim\mathcal{N}(0,\mathbf{I})$, the reverse transition $p_{\theta}(x_{t-1}\mid x_{t})$ is applied iteratively to progressively denoise samples and generate data.
@@ 4 Diffusion Embedded Clustering
Let $x_{t,i}$ denote the noisy version of the $i$-th sample at diffusion timestep $t$, and $x_{0,i}$ its corresponding clean input. Given an unlabeled dataset $\mathcal{D}=\{x_{0,i}\}_{i=1}^{N}$ and the number of clusters $K$, DiEC leverages intermediate activations of a pretrained diffusion model to learn clusterable representations and infer a cluster assignment $\hat{y}_{i}\in\{1,\ldots,K\}$ for each sample.
DiEC first identifies a clustering-friendly location on the diffusion representation trajectory, specified by a layer–timestep pair $(l^{*},t^{*})$. Conditioning on $(l^{*},t^{*})$, it extracts the corresponding features and optimizes a DEC-style clustering objective together with a structural regularizer, yielding stable and well-separated embeddings. Figure 1 provides an overview of the framework, and we present implementation details in the following subsections.
@@ 4.2 Clustering-Aware Optimization
Clustering objective. Following the DEC framework introduced in Sec. 3, we adopt a KL-divergence-based self-training objective on the learned representations $\{z_{i}\}$. Specifically, for each sample $x_{0,i}$, we apply forward noising at the fixed clustering-optimal timestep $t^{*}$ to obtain $x_{t^{*},i}$ and extract its intermediate representation via the CML. To reduce the stochasticity induced by diffusion noise, we draw $M$ independent noise realizations at $t^{*}$, obtain the corresponding representations $\{z_{i}^{(m)}\}_{m=1}^{M}$ through the residual module, and average them to form a stabilized feature
@@
Given $\{z_{i}\}$, we initialize cluster centroids using $k$-means and follow DEC to compute the soft assignments $Q$ and the target distribution $P$. We then minimize the KL divergence $\mathrm{KL}(P\|Q)$ to progressively sharpen high-confidence assignments, and denote this objective as $\mathcal{L}_{\mathrm{KL}}$.
@@ 5 Experiment
Datasets. We evaluate DiEC on four widely used benchmarks for unsupervised clustering: MNIST, USPS, Fashion-MNIST, and CIFAR-10. MNIST and Fashion-MNIST each contain 70,000 grayscale images of resolution $28\times 28$. USPS contains 9,298 grayscale handwritten-digit images with an original resolution of $16\times 16$. CIFAR-10 consists of 60,000 color images of resolution $32\times 32$. All datasets have $10$ ground-truth classes. In our clustering setup, we treat all samples as unlabeled during training and set $K$ to the true number of classes only for evaluation when computing clustering metrics. To match the diffusion model input, all images are normalized as a preprocessing step. The dataset statistics are reported in Table 1.